Jocelyn has test scores of 80, 77 and 89 on her first three Algebra tests. What will she need to score on her 4th test so she will have at least an 85 average?
Answer by John10(297)

Jocelyn has test scores of 80, 77 and 89 on her first three Algebra tests. What will she need to score on her 4th test so she will have at least an 85 average?
Suppose her 4th test grade is x
So the average score is
(80 + 77 + 89 +x)/ 4 at least 85
(246 + x)/4 >= 85
246 + x >= 85*4 = 340
x >= 94
In order to get at 85, she must get 94 on her last test.